C3H8 + O2                       CO2 + H2O

If you start with 14.8g of C3H8 and 3.44g of O2, how many grams of H2O are produced?

Which is the limiting reagent (which reactant created less H2O? _______

How many grams of the non-limiting reagent are actually used?

How many grams of the non-limiting reagent are in excess?




Note : Balance first the equation

It becomes    C3H8 +5O2              3 CO2 + 4H2O



1
Expert's answer
2022-02-22T06:37:02-0500

Molar mass of "C_3H_8=44.1"

Molar mass of O"_2=32.0"

Molar mass of H"_2O=18.0"


Moles of C3H8 "=\frac{14.8}{44.1}=0.3356"


Moles of O"_2=\frac{3.44}{32}=0.1075"


Grams of H2O produced "=0.1075\u00d718"

"=1.935g"


Limiting reagent is O2


Grams of C3H8 used "=0.1075\u00d744.1=4.741g"


Excess of C3H8 "=14.8-4.741"

"=10.059g"